Maria is five times as old as her daughter. Three years ago she was seven times as old as her daughter. How old are they?

Answer by chen(11)

Maria is five times as old as her daughter. Three years ago she was seven times as old as her daughter. How old are they?
Maria’s current age: m
Daughter’s current age: d
Obviously m=5*d
3 years ago, Maria’s age was m-3 and her daughter’s age was d-3, so:
(m-3)=7*(d-3) so
m-3=7d-21. Replacing m from the above relation:
5d-3=7d-21
21-3=7d-5d
2d=18 so d=9 So the daughter is 9 years old and Maria is 45.